In a nuclear fission reaction a heavy nucleus divides to form smaller nuclei and one or more neutrons. Many nuclei can undergo fission, but the fission reactions of uranium-235 and plutonium-239 are the principal ones that generate energy in nuclear power plants. This problem deals with balancing the fission reaction of the uranium-235 isotope as it undergoes bombardment from a neutron.

Part A

When a 235 92U nucleus is bombarded by neutrons (10n) it undergoes a fission reaction, resulting in the formation of two new nuclei and neutrons. The following equation is an example of one such fission process:

235 92U+10n→AZBa+9436Kr+310n

Enter the isotope symbol for the barium (Ba) nucleus in this reaction.

Express your answer as an isotope.

Part B

In another process in which 235 92U undergoes neutron bombardment, the following reaction occurs:

235 92U+10n→AZSr+143 54Xe+310n

Enter the isotope symbol for the strontium (Sr) nucleus in this reaction.

Express your answer as an isotope.
Answer #1

A)

mass balance:

235 + 1= A + 94 + 3

A = 235+1-94-3 = 139

proton number

Ba = 92 + 0 = Z + 36+0

Z = 92-36 = 56

13956Ba

b)

mass balance:

235 + 1 = A + 143 + 3

A = 235+1-143-3

A = 90

proton balance

92 + 0 = Z + 54 + 0

Z = 92-54 = 38

this must be:

9038Sr